OpenStudy (anonymous):

Solve the system of equations involving two circles: x^2+y^2+2x+2y=0 x^2+y^2+4x+6y+12=0

OpenStudy (agreene):

(-2,-2); (-6/5,-12/5) you just need to set them equal to each other and solve for the variables.

OpenStudy (dumbcow):

(x+1)^2 + (y+1)^2 = 2 (x+2)^2 + (y+3)^2 = 1

OpenStudy (anonymous):

(-2,-2); (-6/5,-12/5)